A nationwide survey of college seniors by the University of Michigan revealed that almost 70% disapprove of daily pot smoking. If 12 seniors are selected at random and asked their opinion, find the probability that the number who disapprove of smoking pot daily is greater than 6 and less than 10.

Use 3 significant figures
Answer #1

Here, n = 12, p = 0.7, (1 - p) = 0.3, x1 = 7 and x2 = 9.
As per binomial distribution formula P(X = x) = nCx * p^x * (1 - p)^(n - x)

We need to calculate P(7 <= X <= 9)
P(7 <= X <= 9) = (12C7 * 0.7^7 * 0.3^5) + (12C8 * 0.7^8 * 0.3^4) + (12C9 * 0.7^9 * 0.3^3)
P(7 <= X <= 9) = 0.158 + 0.231 + 0.24
P(7 <= X <= 9) = 0.629